# Explainable AI for Classifying UTI Risk Groups Using a Real-World Linked EHR and Pathology Lab Dataset

### v1

@@ 5.1 Data Preparation
During the 33-month data collection period (from October 1, 2019, to July 1, 2022), multiple potential occurrences of UTIs are identified for some patients, each assigned a likelihood score using the clinician-informed UTI risk estimation framework. As higher likelihood scores reflect greater confidence in the label, to boost the relatively small class sizes for the higher likelihoods, we selected the UTI occurrence with the highest likelihood within each patient’s observation window, favoring the most recent and confident estimate. For each patient, we extracted features from the 12-month period preceding this highest likelihood UTI event, as shown in Figure 1. This standardized observation window across all patients, enabling consistent identification of predictive risk factors influencing the likelihood of UTI occurrence within the next year. Features included in this study are listed in section A5 in Appendix.
@@
To maintain uniformity in temporal analysis, patients with observation windows shorter than 12 months were excluded. Additionally, patients under the age of 18 and those with unknown or unspecified sex were removed to establish a more homogeneous cohort for analysis. The study workflow is described in section A4 in Appendix.
@@ 5.2 Modeling UTI Risk Categories
The clinician-informed UTI risk estimation framework categorizes patients into six distinct likelihood groups. To analyze predictive differences between adjacent risk groups and identify key features predicting transitions between categories, we developed a classification model using XGBoost (Chen and Guestrin 2016) and SHAP (SHapley Additive exPlanations) (Lundberg and Lee 2017) for enhancing explainability.
Pairwise models were constructed to differentiate each set of neighboring risk groups (e.g., 0 vs. 0.2, 0.2 vs. 0.4), with the higher-risk group within each pair considered the positive class. Data for each model was divided into training, validation, and testing sets with the ratio of (3:1:1). The modeling workflow and hyperparameter details for each pairwise model are provided in Appendix A4. These models aimed to identify the most important features distinguishing patients between adjacent UTI risk categories, providing insights about factors associated with an elevated risk of UTI. By leveraging SHAP explainable models, we assessed the impact of individual features on model predictions, aligning machine learning outputs with clinical understanding to ensure interpretability and transparency.
@@ 6.1 Descriptive Analysis
Demographic Information Group
UTI Group [^1] Individuals with a UTI likelihood $>$ 0 Control Group [^2] Individuals with a UTI likelihood $=$ 0
N=147518 N=610602
Age (years)
18-24 9445 (6.4%) 73137 (12.0%)
25-44 49822 (33.8%) 259189 (42.4%)
45-64 37688 (25.5%) 179170 (29.3%)
65-84 39151 (26.5%) 87452 (14.3%)
85+ 11412 (7.7%) 11654 (1.9%)
Gender
Male 45082 (30.6%) 340534 (55.8%)
Female 102436 (69.4%) 270068 (44.2%)
Comorbidities
Incontinent Urinary 538 (0.4%) 405 (0.1%)
Dementia 4693 (3.2%) 3951 (0.6%)
Covid High Risk 22433 (15.2%) 23130 (3.8%)
Covid Increased Risk 84644 (57.4%) 185056 (30.3%)
Organ Transplant 172 (0.1%) 222 (0.0%)
Living Conditions
Housebound 6750 (4.6%) 4031 (0.7%)
Nursing/Caring Home 3304 (2.2%) 2651 (0.4%)
Homeless 159 (0.1%) 502 (0.1%)
@@
An analysis of age and gender distributions reveals a distinct peak among females around the age o 30 within the 0.2 risk group, a pattern not observed in males. This observation may reflect demographic influences on clinical practices. The 0.2 risk group presents a significant proportion, while the 0.4 risk group is small, comprising patients who are dispensed antibiotics sometimes used for UTIs and with negative pathology results or findings involving bacteria rarely associated with UTIs. These risk groups capture cases where clinical suspicion of UTI exists, but uncertainty remains, resulting in a low likelihood risk score. The likelihood of 1.0 group, which corresponds to the highest certainty of UTI diagnosis (e.g., hospitalized patients who had ICD-10 codes confirming UTI), was limited by the relatively small number of severely ill patients in the dataset. From the temporal trend perspective, the 0.2 group shows an obvious increase after mid 2021. This may be attributed to challenges in accessing healthcare during the pandemic and a subsequent rise in clinical activities as restrictions were eased. Interestingly, a slight upward trend in the 0.6 and 0.8 groups after April 2022 is discovered, possibly reflecting the impact of policy changes or adjustments in clinical practices.
@@ 6.2 Model Performance and Feature Importance
To differentiate patients between neighbouring UTI likelihood categories, pairwise models were constructed for adjacent groups (e.g., 0 vs. 0.2, 0.2 vs. 0.4). Table 2 summarizes the performance metrics of these models over the testing set, including accuracy, precision, recall, F1-score, and AUC-ROC. Figure 6 displays the ROC curves for all the pairwise models, providing a visual comparison of their discriminative ability across the adjacent UTI risk categories.
@@
0.0 vs 0.2 0.90 0.77 0.28 0.41 0.80
0.2 vs 0.4 0.98 $<$0.01 $<$0.01 $<$0.01 0.62
0.4 vs 0.6 0.95 0.95 1.00 0.98 0.78
0.6 vs 0.8 0.74 0.74 0.40 0.52 0.73
0.8 vs 1.0 0.96 0.73 0.46 0.56 0.98
Table 2: Performance metrics of the pairwise models.
Figure 6: Roc-curve.
Figure 7: Top 15 features identified by SHAP for the five pairwise XGBoost models distinguishing adjacent UTI risk categories.
The models demonstrate varying effectiveness in distinguishing between adjacent UTI risk groups. The 0.0 vs 0.2 model achieves high accuracy (0.90) and an AUC-ROC of 0.80, but its low recall suggests difficulty in identifying true positives with subtle feature differences. The 0.2 vs 0.4 model has high accuracy but performs poorly in other metrics, likely due to extreme class imbalance. In contrast, the 0.4 vs 0.6 model performs relatively well in distinguishing each other across all metrics. This result aligns with human expectations of risk categorization, as the natural threshold at 0.5 intuitively separates patients with higher likelihood scores above 0.5 from those below 0.5. However, the AUC-ROC of 0.78 still indicates some challenges in distinguishing between these adjacent risk groups. The 0.6 vs 0.8 model shows moderate performance (AUC-ROC 0.73) but the recall (0.40) and F1-score (0.52) suggest that the model has limitations in detecting instances of the higher-risk group. Finally, the 0.8 vs 1.0 model achieves excellent discrimination with an AUC-ROC of 0.98, though its lower recall and precision reflects some misclassification, likely due to class imbalance of the positive class in this model.
The SHAP summary plots (Figure 7) provide insights for each pairwise comparison of adjacent UTI risk categories. Key patterns and trends were observed, reflecting the evolving importance of clinical and demographic features across the UTI likelihood groups. Features such as the number of previous antibiotics dispensed and the number of urine specimens processed are critical in distinguishing low-likelihood risk groups, highlighting the relevance of prior clinical interventions in shaping risk predictions for less certain UTI cases. In the 0.4 vs 0.6 model, the focus shifted towards more specific clinical features, such as the number of UTI-specific antibiotics dispensed. This model also highlighted the growing importance of comorbidities, including dementia and prior resistance discoveries, which became more significant in distinguishing the intermediate risk groups. For the 0.6 vs 0.8 and 0.8 vs 1.0 models, age displayed a clearer trend, with older patients more likely to be classified in the higher-risk groups. Features such as prior resistance discoveries and the number of urine specimens processed were dominant in the model of the 0.6 vs 0.8. The model that differentiates patients with hospital-confirmed UTI diagnoses, further highlighted severe case predictors such as the number of blood specimens processed, COVID-19 risk histories, housebound status, dementia.
These findings demonstrate the progressive shift in feature importance, from general clinical and demographic variables in lower-risk comparisons to more specific and severe clinical factors in higher-risk groups. The clearer trend for age in higher-risk categories aligns with clinical expectations, reinforcing the interpretability of the models and their alignment with the clinical expectations.
@@ 7 Discussion
In this paper, we introduced a unique linked EHR dataset and detailed our pre-processing and curation methods tailored for UTI prediction. This enriched dataset integrates information from multiple healthcare settings, offering potential for advancing AI-driven healthcare solutions. However, it also presents several limitations. For instance, the dispensation data we acquired differs from the prescribing records commonly used in similar studies. Dispensation data records medications dispensed to patients rather than prescribed, providing a more accurate representation of actual medication usage. Nevertheless, it does not confirm whether patients completed the course or initialized it at all. Additionally, while urine sample processing data is included in the dataset, discrepancies may arise because the processing date may not align with the date the sample was collected, introducing potential temporal inconsistencies. Our initial clinician-guided framework for data annotation facilitates supervised learning and reflects how clinicians assess UTI risk based on three primary variables. However, this framework is based on current clinical practice and depends on a restricted set of clinical indicators, potentially overlooking subtle or unrecognized factors influencing UTI risk. Moreover, the 0.4 risk group contains relatively few cases, which may affect the robustness of the analysis. To address this, merging it with other groups, guided by AI to explore possible combinations, could potentially enhance reliability. Addressing the challenges posed by incomplete data and the lack of explicit UTI labels remains critical. These limitations introduce uncertainty into the analysis, emphasizing the importance of robust estimation methods to enhance the reliability of risk assessment. Future work will aim to refine this approach by transitioning to explainable AI-based predictive models capable of uncovering latest risk groups as well as broader patterns, providing greater insights into predictive factors and modifiable risk factors.
Our study highlights the critical role of integrating data curation, risk estimation frameworks, and interpretable AI techniques to improve the real-world applicability of predictive models in healthcare. By leveraging a diverse and enriched dataset, we aim to advance not only the accuracy of predictive analytics but also their equity and effectiveness, ultimately contributing to better healthcare solutions.
Appendix A Appendices
For additional details and information, please refer to the appendix file available at the following link: Appendix
Acknowledgements
This project is supported by funding from Engineering and Physical Sciences Research Council (EPSRC). The authors thank Debbie Campbell and Elizabeth Jonas for their time and effort in reviewing.

### v2

@@ 5.1 Data Preparation
During the 33-month data collection period (from October 1, 2019, to July 1, 2022), multiple potential occurrences of UTIs are identified for some patients, each assigned a likelihood score using the clinician-informed UTI risk estimation framework. As higher likelihood scores reflect greater confidence in the label, to boost the relatively small class sizes for the higher likelihoods, we selected the UTI occurrence with the highest likelihood within each patient’s observation window, favoring the most recent and confident estimate. For each patient, we extracted features from the 12-month period preceding this highest likelihood UTI event, as shown in Figure 1. This standardized observation window across all patients, enabling consistent identification of predictive risk factors influencing the likelihood of UTI occurrence within the next year. Features included in this study are listed in Appendix A5.
@@
To maintain uniformity in temporal analysis, patients with observation windows shorter than 12 months were excluded. Additionally, patients under the age of 18 and those with unknown or unspecified sex were removed to establish a more homogeneous cohort for analysis. The study workflow is described in Appendix A4.
@@ 5.2 Modeling UTI Risk Categories
The clinician-informed UTI risk estimation framework categorizes patients into six distinct likelihood groups. To analyze predictive differences between adjacent risk groups and identify key features predicting transitions between categories, we developed pairwise classification models using XGBoost xgboost and SHAP (SHapley Additive exPlanations) SHAP for enhancing explainability. XGBoost was selected based on its superior predictive performance compared to logistic regression and random forest models in prior experiments. SHAP was chosen for its strong theoretical foundation based on Shapley values and its robustness and consistency across models, ensuring reliable insights into feature contributions and interactions. Each pairwise models were designed to differentiate neighboring risk groups (e.g., 0 vs. 0.2, 0.2 vs. 0.4), with the higher-risk group within each pair considered the positive class. Data for each model was divided into training, validation, and testing sets with the ratio of (3:1:1) with the modeling workflow and hyperparameter details provided in Appendix A4. As the 0.4 and 1 UTI groups are relatively small, there is potential of class imbalance when modeling adjacent classes. To mitigate this issue, we experimented with three approaches: (1) no intervention, (2) applying the scale_pos_weight hyperparameter in XGBoost, and (3) employing SMOTE (Synthetic Minority Over-sampling Technique) SMOTE for oversampling the minority class. While all three methods showed consistent Shapley feature explanations, the weighted balancing (using scale_pos_weight) was ultimately chosen as in our study. This method offered a good balance between model performance and stability in feature importance, without introducing the alterations in feature distributions observed with SMOTE. The pairwise models identify the important features distinguishing patients between adjacent UTI risk categories, providing insights about factors associated with elevated UTI risk. By leveraging SHAP explainable models to assess the impact of individual features on model predictions, we aim to align machine learning outputs with clinical understanding to ensure interpretability and transparency.
@@ 6.1 Descriptive Analysis
Demographics Groups
Control Group [^1] Individuals with a UTI likelihood $=$ 0 UTI Group [^2] Individuals with a UTI likelihood $>$ 0
Overall Overall 0.2 0.4 0.6 0.8 1.0
N = 610602 147518 84280 1969 38278 21539 1452
Age (years) (%)
18-24 12.0 6.4 5.8 5.9 9.0 4.5 1.2
25-44 42.4 33.8 38.9 37.0 31.2 19.6 5.9
45-64 29.3 25.5 25.8 26.5 26.4 23.6 14.2
65-84 14.3 26.5 23.5 24.3 25.0 39.9 45.0
85+ 1.9 7.7 5.9 6.3 8.3 12.4 33.6
Gender (%)
Male 55.8 30.6 37.8 36.4 19.1 20.6 51.4
Female 44.2 69.4 62.2 63.6 80.9 79.4 48.6
Comorbidities (%)
Incontinent Urinary 0.1 0.4 0.2 0.3 0.5 0.7 0.9
Dementia 0.6 3.2 2.0 1.7 4.7 4.6 12.8
Covid High Risk 3.8 15.2 15.2 16.6 13.3 16.9 38.8
Covid Increased Risk 30.3 57.4 55.0 56.6 54.8 69.2 91.9
Organ Transplant $<$0.1 0.1 0.1 0.4 0.1 0.1 0.3
Living Conditions (%)
Housebound 0.7 4.6 3.4 5.1 5.3 6.8 22.0
Nursing/Caring Home 0.4 2.2 1.3 1.9 3.3 3.7 7.1
Homeless 0.1 0.1 0.1 0.2 0.1 $<$0.1 0.1
@@
An analysis of age and gender distributions reveals a distinct peak among females around the age of 30 within the 0.2 risk group, a pattern not observed in males. This observation may reflect demographic influences on clinical practices, such as urine samples collected during pregnancy, where treatment is provided only if symptomatic or a pathogen is identified. The 0.2 risk group presents a significant proportion, while the 0.4 risk group is small, comprising patients who are dispensed antibiotics sometimes used for UTIs and with negative pathology results or findings involving bacteria rarely associated with UTIs. These risk groups capture cases where clinical suspicion of UTI exists, but uncertainty remains, resulting in a low likelihood risk score. The likelihood of 1.0 group, which corresponds to the highest certainty of UTI diagnosis (e.g., hospitalized patients who had ICD-10 codes confirming UTI), was limited by the relatively small number of severely ill patients in the dataset. From the temporal trend perspective, the 0.2 group shows an obvious increase after mid 2021. This may be attributed to challenges in accessing healthcare during the pandemic and a subsequent rise in clinical activities as restrictions were eased. Interestingly, a slight upward trend in the 0.6 and 0.8 groups after April 2022 is discovered, possibly reflecting the impact of policy changes or adjustments in clinical practices.
t]
@@ 6.2 Model Performance and Feature Importance
To differentiate patients between neighbouring UTI likelihood categories, pairwise models were constructed for adjacent groups (e.g., 0 vs. 0.2, 0.2 vs. 0.4). However, from the clinical perspective underlying the UTI risk framework, the 0.4 risk group is notably small, with only one instance allocated to this category, making it challenging to model effectively. To address this, we constructed an additional pairwise model for the 0.2 vs. 0.6 groups to explore the differences between these two categories. Table 6.2 summarizes the performance metrics of these models over the testing set, including accuracy, precision, recall, F1-score, and AUC-ROC. Figure 6 displays the ROC curves for all the pairwise models, providing a visual comparison of their discriminative ability across the adjacent UTI risk categories. The models demonstrate varying effectiveness in distinguishing between adjacent UTI risk groups. The model differentiating higher-risk groups 0.8 vs. 1.0 generally performs better, with high accuracy (0.94) and AUC-ROC (0.98), suggesting that these groups are more easily distinguishable. However, models distinguishing between lower-risk groups, like 0.0 vs. 0.2 and 0.2 vs. 0.4, tend to show lower precision and F1-scores, which may be influenced by class imbalance, particularly the small size of the 0.4 group. The 0.2 vs. 0.6 model, introduced to explore the differentiation between a low and moderately high-risk group, shows moderate accuracy (0.71) but a relatively low F1-score (0.58), reflecting the challenge of distinguishing between groups that are distinct but still share similar characteristics.
Table 2: Performance metrics of the pairwise models.
@@
[…]
0.2 vs 0.4 0.87 0.06 0.33 0.10 0.62
0.4 vs 0.6 0.69 0.98 0.69 0.81 0.78
0.6 vs 0.8 0.72 0.62 0.53 0.57 0.73
0.8 vs 1.0 0.94 0.51 0.94 0.66 0.98
0.2 vs 0.6 0.71 0.53 0.64 0.58 0.77
t]
Figure 6: Roc-curve for all the pairwise models over the testing data.
The SHAP summary plots (Figure 7) reveal evolving patterns in feature importance as the risk groups progress, reflecting both clinical and demographic factors that influence UTI likelihood. In the 0.0 vs. 0.2 model, features related to prior clinical interventions, such as the number of previous antibiotics dispensed and the number of urine specimens processed, play a prominent role in distinguishing between low-risk groups, suggesting that historical clinical events within one year contribute to predicting risk in these less certain cases. In the 0.4 vs. 0.6 model, the focus shifted towards more specific clinical features, including the number of UTI-specific antibiotics dispensed. Additionally, comorbidities like dementia and prior resistance discoveries become more relevant, indicating their growing importance in identifying patients at moderate risk. In the 0.2 vs. 0.6 model, which differentiates between low and moderately high-risk patients, the number of UTI-specific antibiotics dispensed again stands out as a key feature. For the 0.6 vs. 0.8 and 0.8 vs 1.0 models, age displays a clearer trend, with older patients more likely to be classified into the higher-risk groups. Features such as prior resistance discoveries and the number of urine specimens processed are particularly influential in the model of the 0.6 vs. 0.8. When distinguishing patients in the group of 0.8 vs. 1.0, the model highlights severe case predictors, such as housebound status and the number of previous hospital visits, underscoring the role of these factors in patients with hospital-confirmed UTI diagnoses (UTI likelihood $=$ 1.0).
Figure 7: Top 15 features identified by SHAP for the six pairwise XGBoost models distinguishing adjacent UTI risk categories.
The results from SHAP help understand how the models prioritize different features at each stage of the risk estimation process. These observations suggest that the importance of various clinical and demographic features changes across risk categories, with certain factors becoming more prominent as risk increases. While these results provide useful insights, future work will be necessary to refine the categorization of risk groups, particularly by adjusting the groupings to better capture the variability in patient risk profiles and improve model robustness.
@@ 7 Discussion
In this study, we introduced a unique linked EHR dataset and detailed the pre-processing and curation methods tailored for UTI prediction. Despite its potential to advance AI-driven healthcare solutions, the dataset has several limitations. The use of dispensation data, rather than prescribing records commonly used in similar studies, provides a more accurate representation of actual medication usage but lacks confirmation of whether patients initiated or completed their medication courses. Additionally, urine sample processing data may have temporal inconsistencies, as the processing date does not always align with the sample collection date. Our initial clinician-guided framework, while grounded in current clinical practices, focuses on a limited set of clinical indicators and may overlook subtle or unrecognized factors that influence UTI risk. Furthermore, the relatively small size of the 0.4 risk group may affect the robustness of analyses, suggesting the need to consider merging it with adjacent groups or exploring alternative groupings using AI-driven methods. Finally, addressing the challenges posed by incomplete data and the lack of explicit UTI labels remains critical. These limitations introduce uncertainty into the analysis, emphasizing the importance of robust estimation methods to enhance the reliability of risk assessment.
To address these limitations, future work will focus on refining the risk estimation framework through advanced sub-grouping and clustering techniques, guided by explainable AI, to uncover more comprehensive patterns in the data and enhance UTI risk group definitions. This includes exploring data-driven strategies for reclassifying smaller or less robust groups, such as the 0.4 risk category, to improve analytical reliability. Additionally, efforts will be directed toward investigating subgroup-specific risk factors to ensure fairness and inclusivity, particularly for underrepresented patient populations. Incorporating novel clinical indicators and integrating longitudinal patient histories may provide deeper insights into the predictors of UTI risk. Enhanced validation through external datasets and collaborations with diverse healthcare settings will further ensure the robustness and generalizability of the findings.
8 Conclusion
This study represents a first step in developing a clinician-guided framework for UTI risk predictors investigation with explainable AI. While preliminary, our results offer insights into the predictors of UTI risk and demonstrate the importance of integrating data curation, risk estimation frameworks, and interpretable AI techniques. However, further validation and refinement are needed before these models can be implemented in clinical practice. By leveraging a diverse and enriched dataset, we aim to improve the reliability of predictive analytics, promote healthcare equity, and enhance clinical decision-making and patient outcomes.
See pages - of Appendix.pdf